OpenStudy (anonymous):

Let |4 0 1| A= |2 3 2| |25 0 4| . For each eigenvalue λ, find the rank of the matrix λI - A.

OpenStudy (anonymous):

@phi @ganeshie8

OpenStudy (phi):

replace the diagonal 4 3 4 with (4-L), 3-L , 4-L then expand it as if you were finding the determinant. This gives you the "characteristic equation" that you solve to find L

OpenStudy (anonymous):

Thanks. Uhm, how about the rank?

OpenStudy (phi):

for each eigenvalue, you find the matrix λI - A now do row reduction and count the number of pivots

OpenStudy (phi):

refresh your browser if it does not display properly

OpenStudy (anonymous):

Thank youuu :D

OpenStudy (phi):

if you expand using the middle column (it has 2 zeros, so it's convenient), the characteristic equation is (3-x)[ (4-x)^2 -25] = 0 or (3-x)[ x^2 -8x +16 -25]= 0 (x-3) (x^2 -8x -9) = 0 we see that this will be zero if (x-3)= 0 or x^2 -8x-9 = 0